# The fractional $p$-Laplacian on hyperbolic spaces

### v2

@@ Abstract.
We present three equivalent definitions of the fractional $p$-Laplacian $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}$, $10<s<$ 1, $p>1$, with normalizing constants, on hyperbolic spaces. The explicit values of the constants enable us to study the convergence of the fractional $p$-Laplacian to the $p$-Laplacian as $s\to 1^{-}$.
@@ 1. Introduction
The aim of this paper is two-fold. First, we extend representation formulas in [2] to the nonlinear regime on hyperbolic spaces. Specifically, we define the fractional $p$-Laplacian $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}$ for $n\in\mathbb{N}$, $10<s<$ 1, and $p>1$ and provide two additional equivalent definitions via the heat semigroup and the Caffarelli–Silvestre extension. Note that a definition based on the Fourier transform is not available due to the nonlinearity of the operator. Second, we investigate the pointwise convergence of $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}u(x)$ as $s\to 1^{-}$. For this purpose, the values of the normalizing constants in these definitions are provided explicitly.
@@
Let $n\geq 2$, $10<s<$ 1, and $p>1$. The fractional $p$-Laplacian on $\mathbb{H}^{n}$ is defined by
@@
$c_{n,s,p}=\frac{p}{2}\frac{\sqrt{\pi}/2}{\Gamma(\frac{p+1}{2})}\frac{2^{2s}\Gamma(\frac{n+sp}{2})}{\pi^{\frac{n}{2}}|\Gamma(-s)|},\quad C_{1}=\frac{1}{2^{\frac{n-2+sp}{2}}\Gamma(\frac{n+sp}{2})}\left(\frac{n-1}{2}\right)^{\frac{1+sp}{2}},$
@@
Note that $(-\Delta_{\mathbb{H}^{n}})_{p}^{s}u(x)$ is well defined for $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$ (assuming also that $\nabla u(x)\neq 0$ if $p\in(1,\frac{2}{2-s}]$), where $C^{2}_{b}(\mathbb{H}^{n})$ denotes the space of bounded $C^{2}$-functions on $\mathbb{H}^{n}$. We refer the reader to [28, Section 3], where the authors provide the well-definedness of $(-\Delta_{\mathbb{R}^{n}})_{p}^{s}u(x)$; the same proof also works on hyperbolic spaces.
@@
The singular kernel $\rho^{-n-sp}$ of the fractional $p$-Laplacian on Euclidean space $\mathbb{R}^{n}$ is homogeneous of degree $-n-sp$. This is a natural consequence of the invariance of the scale of the operator. However, such homogeneity cannot be expected in the hyperbolic setting because hyperbolic geometry comes into play. In fact, we prove the following proposition. (In what follows, we write $A(\rho)\sim B(\rho)$ as $\rho\to 0^{+}$ (resp. $\rho=\infty$) to mean that there exist constants $c,C>0$ and $\rho_{0}>0$ such that $cB(\rho)\leq A(\rho)\leq CB(\rho)$ for all $\rho\in(0,\rho_{0})$ (resp. $\rho\in(\rho_{0},\infty)$).)
@@
Let $n\in\mathbb{N}$, $10<s<$ 1, and $p>1$. Let $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$. If $p\in(1,\frac{2}{2-s}]$, assume in addition that $\nabla u(x)\neq 0$. The fractional $p$-Laplacian on $\mathbb{H}^{n}$ is defined by
@@
Let $n\in\mathbb{N}$, $10<s<$ 1, and $p>1$. Let $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$. If $p\in(1,\frac{2}{2-s}]$, assume $\nabla u(x)\neq 0$ additionally. Then
@@ 2.1. The hyperbolic space
with the Lorentzian metric $-\mathrm{d}x_{0}^{2}+\mathrm{d}x_{1}^{2}+\cdots+\mathrm{d}x_{n}^{2}$ in $\mathbb{R}^{n+1}$. The Lorentzian metric induces the natural internal product
$[x,\xi]=x_{0}\xi_{0}-x_{1}\xi_{1}-\cdots-x_{n}\xi_{n}$
@@ 2.2. The modified Bessel function
(2.5) $-\partial_{\rho}(\mathscr{K}_{\nu,a}(f(\rho)))=af^{\prime}(\rho)f(\rho)\mathscr{K}_{\nu+1,a}(f(\rho))$
@@ 3. Nonlinear Bochner’s formula
The main goal of this section is to prove Theorem 1.3. One of the crucial tools in the proof is the explicit formula for the heat kernel on hyperbolic spaces, which is given in [23] as follows:
(3.1) $p(t,\rho)=\frac{1}{(2\pi)^{m}}\frac{1}{(4\pi t)^{1/2}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}e^{-m^{2}t-\frac{\rho^{2}}{4t}}$
@@
(3.2) $p(t,\rho)=\frac{1}{2(2\pi)^{m+1/2}}t^{-3/2}e^{-\frac{(2m-1)^{2}}{4}t}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{re^{-\frac{r^{2}}{4t}}}{\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r$
@@
$w(x,t)=\int_{\mathbb{H}^{n}}p(t,d(x,\xi))f(\xi)\,\mathrm{d}\xi,$
provided that $f$ is a bounded continuous function.
@@
$F_{m}(r):=\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right),$
where $\mathscr{K}_{\nu,a}$ is the function given in (2.4). Then, $F_{m}$ is integrable on $(\rho,\infty)$ and satisfies
@@
$-\partial_{r}\left(\frac{(e^{r}+e^{-r})^{j-1}}{(e^{r}-e^{-{r}})^{j}}\right)=j\frac{(e^{r}+e^{-r})^{j}}{(e^{r}-e^{-r})^{j+1}}-(j-1)\frac{(e^{r}+e^{-r})^{j-2}}{(e^{r}-e^{-r})^{j-1}}.$
@@
$\begin{split}\int_{\rho}^{\infty}F_{m}(r)\,\mathrm{d}r&=\int_{\rho}^{\infty}2\partial_{r}\left(\sqrt{\cosh r-\cosh\rho}\right)\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right)\mathrm{d}r\\ &=\int_{\rho}^{\infty}2\sinh r\sqrt{\cosh r-\cosh\rho}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\mathscr{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right)\mathrm{d}r.\end{split}$
@@
$\rho\mapsto\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\nu,a}(\rho)$
@@
$\mathscr{K}_{\nu,a}(\rho)=\frac{a^{\nu}}{2^{\nu+1}}\int_{0}^{\infty}e^{-t-\frac{(a\rho)^{2}}{4t}}t^{-\nu-1}\,\mathrm{d}t=\frac{1}{2(2a)^{\nu}}\int_{0}^{\infty}\frac{1}{t^{1/2}}e^{-a^{2}t-\frac{\rho^{2}}{4t}}\frac{\mathrm{d}t}{t^{\nu+1/2}}.$
@@
$\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\nu,a}(\rho)=\int_{0}^{\infty}e^{(m^{2}-a^{2})t}p(t,\rho)\frac{\mathrm{d}t}{t^{\nu+1/2}}.$
The conclusion follows from the positivity of the heat kernel $p$. ∎
@@
Let $n\in\mathbb{N}$, $10<s<$ 1, and $p>1$. The kernel $\mathcal{K}_{n,s,p}$ is positive.
@@
$K(\rho)=\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\mathscr{K}_{\nu,a}(\rho)$
@@
$K(\rho)=\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\mathscr{K}_{\nu,a}(r)\mathrm{d}r$
@@
[…]
@@
$K(\rho)$ $=\int_{\rho}^{\infty}2\sinh r\sqrt{\cosh r-\cosh\rho}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n+2}{2}}\mathscr{K}_{\nu,a}(r)\,\mathrm{d}r$
@@
(3.8) $\int_{0}^{\infty}e^{-m^{2}t-\frac{\rho^{2}}{4t}}t^{-\frac{3+sp}{2}}\,\mathrm{d}t=m^{1+sp}\int_{0}^{\infty}e^{-t-\frac{(m\rho)^{2}}{4t}}t^{-\frac{3+sp}{2}}\,\mathrm{d}t=2(2m)^{\frac{1+sp}{2}}\mathscr{K}_{\frac{1+sp}{2},m}(\rho).$
@@
$\begin{split}\int_{0}^{\infty}e^{-\frac{(2m-1)^{2}}{4}t-\frac{r^{2}}{4t}}t^{-\frac{5+sp}{2}}\,\mathrm{d}t&=2(2m-1)^{\frac{3+sp}{2}}\mathscr{K}_{\frac{3+sp}{2},\frac{2m-1}{2}}(r)\\ &=4(2m-1)^{\frac{1+sp}{2}}\left(\frac{-\partial_{r}}{r}\right)\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r).\end{split}$
@@
$\begin{split}&C_{2}\int_{0}^{\infty}e^{t\Delta_{\mathbb{H}^{n}}}[g_{\varepsilon}(x,\cdot)](x)\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}\\ &=c_{n,s,p}C_{1}\int_{\mathbb{H}^{n}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{(-\partial_{r})\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r\,g_{\varepsilon}(x,\xi)\,\mathrm{d}\xi\\ &=c_{n,s,p}C_{1}\int_{\mathbb{H}^{n}}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,g_{\varepsilon}(x,\xi)\,\mathrm{d}\xi,\end{split}$
where we used Lemma 3.1 $(m-1)$-times with $\nu=\frac{1+sp}{2}$, $a=\frac{2m-1}{2}$, and $y=0$ in the last equality. This proves (3.4) in the even-dimensional case.

### v3

@@ Abstract.
We present three equivalent definitions of the fractional $p$-Laplacian $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}$, $0<s<1$, $p>1$, with normalizing constants, on hyperbolic spaces. The explicit values of the constants enable us to study the convergence of the fractional $p$-Laplacian to the $p$-Laplacian as $s\to 1^{-}$.
@@ 1. Introduction
The aim of this paper is two-fold. First, we extend representation formulas in [2] to the nonlinear regime on hyperbolic spaces. Specifically, we define the fractional $p$-Laplacian $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}$ for $n\in\mathbb{N}$, $0<s<1$, and $p>1$ and provide two additional equivalent definitions via the heat semigroup and the Caffarelli–Silvestre extension. Note that a definition based on the Fourier transform is not available due to the nonlinearity of the operator. Second, we investigate the pointwise convergence of $(-\Delta_{\mathbb{H}^{n}})^{s}_{p}u(x)$ as $s\to 1^{-}$. For this purpose, the values of the normalizing constants in these definitions are provided explicitly.
@@
Let $n\geq 2$, $0<s<1$, and $p>1$. The fractional $p$-Laplacian on $\mathbb{H}^{n}$ is defined by
@@
(1.2) $c_{n,s,p}=\frac{p}{2}\frac{\sqrt{\pi}/2}{\Gamma(\frac{p+1}{2})}\frac{2^{2s}\Gamma(\frac{n+sp}{2})}{\pi^{\frac{n}{2}}|\Gamma(-s)|},\quad C_{1}=\frac{1}{2^{\frac{n-2+sp}{2}}\Gamma(\frac{n+sp}{2})}\left(\frac{n-1}{2}\right)^{\frac{1+sp}{2}},$
@@
The singular kernel $\rho^{-n-sp}$ of the fractional $p$-Laplacian on Euclidean space $\mathbb{R}^{n}$ is homogeneous of degree nsp-n-sp. This is a natural consequence of the invariance of the scale of the operator. However, such homogeneity cannot be expected in the hyperbolic setting because hyperbolic geometry comes into play. In fact, Proposition 1.2 below exhibits the behavior of the kernel $\mathcal{K}_{n,s,p}$.
Here and in what follows, we write $A(\rho)\sim B(\rho)$ as $\rho\to 0^{+}$ (resp. $\rho=\infty$) to mean that there exist constants $c,C>0$ and $\rho_{0}>0$ such that $cB(\rho)\leq A(\rho)\leq CB(\rho)$ for all $\rho\in(0,\rho_{0})$ (resp. $\rho\in(\rho_{0},\infty)$).
@@
We remark that the well-definedness of $(-\Delta_{\mathbb{H}^{n}})_{p}^{s}u(x)$ for $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$ (assuming also that $\nabla u(x)\neq 0$ if $p\in(1,\frac{2}{2-s}]$), where $C^{2}_{b}(\mathbb{H}^{n})$ denotes the space of bounded $C^{2}$-functions on $\mathbb{H}^{n}$, can be checked by using Proposition 1.2. Indeed, the proof of [28, Lemma 3.6] works on hyperbolic spaces in the exact same way, once we choose sufficiently small $\varepsilon>0$ so that $\mathcal{K}_{n,s,p}(d(x,\xi))\sim d(x,\xi)^{-n-sp}$ in the ball $B_{\varepsilon}(x)=\{\xi\in\mathbb{H}^{n}:d(x,\xi)<\varepsilon\}$ with the help of Proposition 1.2. This proves the finiteness of the integral in (1.1) over $B_{\varepsilon}(x)$. On the other hand, by using Proposition 1.2 again, we find a large $R>0$ so that $\mathcal{K}_{n,s,p}(d(x,\xi))\sim d(x,\xi)^{-1-\frac{sp}{2}}e^{-(n-1)d(x,\xi)}$ outside $B_{R}(x)$. This proves that the integral in (1.1) over $\mathbb{H}^{n}\setminus B_{R}(x)$ is estimated by
$\max\{2^{p-2},1\}\|u\|_{L^{\infty}(\mathbb{H}^{n})}^{p-1}\int_{\mathbb{S}^{n-1}}\int_{R}^{\infty}\rho^{-1-\frac{sp}{2}}e^{-(n-1)\rho}\sinh^{n-1}\rho\,\mathrm{d}\rho\,\mathrm{d}\omega,$
which is finite. The finiteness of the remaining integral over $B_{R}(x)\setminus B_{\varepsilon}(x)$ is obvious since the integrand is bounded in a bounded region.
@@
Let $n\in\mathbb{N}$, $0<s<1$, and $p>1$. Let $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$. If $p\in(1,\frac{2}{2-s}]$, assume in addition that $\nabla u(x)\neq 0$. The fractional $p$-Laplacian on $\mathbb{H}^{n}$ is defined by
@@
Let $n\in\mathbb{N}$, $0<s<1$, and $p>1$. Let $u\in C^{2}_{b}(\mathbb{H}^{n})$ and $x\in\mathbb{H}^{n}$. If $p\in(1,\frac{2}{2-s}]$, assume $\nabla u(x)\neq 0$ additionally. Then
@@
The authors would like to thank the anonymous referees for their careful reading of the manuscript and for providing helpful comments.
@@ 2.1. The hyperbolic space
with the Lorentzian metric $\mathrm{d}x_{0}^{2}+\mathrm{d}x_{1}^{2}+\cdots+\mathrm{d}x_{n}^{2}$ in $\mathbb{R}^{n+1}$. The Lorentzian metric induces the natural internal product
$x,\xi]=x_{0}\xi_{0}-x_{1}\xi_{1}-\cdots-x_{n}\xi_{n}$
@@ 2.2. The modified Bessel function
(2.5) $\partial_{\rho}(\widetilde{K}_{\nu,a}(f(\rho)))=af^{\prime}(\rho)f(\rho)\widetilde{K}_{\nu+1,a}(f(\rho))$
@@ 3. Nonlinear Bochner’s formula
The main goal of this section is to prove Theorem 1.3. One of the crucial tools in the proof is the explicit formula for the heat kernel $h$ on hyperbolic spaces, which is given in [23] as follows:
(3.1) $h(t,\rho)=\frac{1}{(2\pi)^{m}}\frac{1}{(4\pi t)^{1/2}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}e^{-m^{2}t-\frac{\rho^{2}}{4t}}$
@@
(3.2) $h(t,\rho)=\frac{1}{2(2\pi)^{m+1/2}}t^{-3/2}e^{-\frac{(2m-1)^{2}}{4}t}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{re^{-\frac{r^{2}}{4t}}}{\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r$
@@
(3.3) $w(x,t)=\int_{\mathbb{H}^{n}}h(t,d(x,\xi))f(\xi)\,\mathrm{d}\xi,$
provided that $f$ is a bounded continuous function. In other words, the heat semigroup $e^{t\Delta_{\mathbb{H}^{n}}}[f]$ is given by (3.3).
@@
$F_{m}(r):=\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right),$
where $\widetilde{K}_{\nu,a}$ is the function given in (2.4). Then, $F_{m}$ is integrable on $(\rho,\infty)$ and satisfies
@@
$\partial_{r}\left(\frac{(e^{r}+e^{-r})^{j-1}}{(e^{r}-e^{-{r}})^{j}}\right)=j\frac{(e^{r}+e^{-r})^{j}}{(e^{r}-e^{-r})^{j+1}}-(j-1)\frac{(e^{r}+e^{-r})^{j-2}}{(e^{r}-e^{-r})^{j-1}}.$
@@
$\begin{split}\int_{\rho}^{\infty}F_{m}(r)\,\mathrm{d}r&=\int_{\rho}^{\infty}2\partial_{r}\left(\sqrt{\cosh r-\cosh\rho}\right)\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right)\mathrm{d}r\\ &=\int_{\rho}^{\infty}2\sinh r\sqrt{\cosh r-\cosh\rho}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\widetilde{K}_{\nu,a}\left(\sqrt{r^{2}+y^{2}}\right)\mathrm{d}r.\end{split}$
@@
$\rho\mapsto\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\nu,a}(\rho)$
@@
$\widetilde{K}_{\nu,a}(\rho)=\frac{a^{\nu}}{2^{\nu+1}}\int_{0}^{\infty}e^{-t-\frac{(a\rho)^{2}}{4t}}t^{-\nu-1}\,\mathrm{d}t=\frac{1}{2(2a)^{\nu}}\int_{0}^{\infty}\frac{1}{t^{1/2}}e^{-a^{2}t-\frac{\rho^{2}}{4t}}\frac{\mathrm{d}t}{t^{\nu+1/2}}.$
@@
$\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\nu,a}(\rho)=\int_{0}^{\infty}e^{(m^{2}-a^{2})t}h(t,\rho)\frac{\mathrm{d}t}{t^{\nu+1/2}}.$
The conclusion follows from the positivity of the heat kernel $h$. ∎
@@
Let $n\in\mathbb{N}$, $0<s<1$, and $p>1$. The kernel $\mathcal{K}_{n,s,p}$ is positive.
@@
$K(\rho)=\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\widetilde{K}_{\nu,a}(\rho)$
@@
$K(\rho)=\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\widetilde{K}_{\nu,a}(r)\mathrm{d}r$
@@
$\leq C\rho^{-\frac{1}{2}-\nu}e^{-(a+\frac{2m-1}{2})\rho}\int_{0}^{\infty}\frac{1}{\sqrt{\sinh\frac{t}{2}}}dt$
@@
$K(\rho)$ $=\int_{\rho}^{\infty}2\sinh r\sqrt{\cosh r-\cosh\rho}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n+2}{2}}\widetilde{K}_{\nu,a}(r)\,\mathrm{d}r$
@@
(3.9) $\int_{0}^{\infty}e^{-m^{2}t-\frac{\rho^{2}}{4t}}t^{-\frac{3+sp}{2}}\,\mathrm{d}t=m^{1+sp}\int_{0}^{\infty}e^{-t-\frac{(m\rho)^{2}}{4t}}t^{-\frac{3+sp}{2}}\,\mathrm{d}t=2(2m)^{\frac{1+sp}{2}}\widetilde{K}_{\frac{1+sp}{2},m}(\rho).$
@@
$\begin{split}\int_{0}^{\infty}e^{-\frac{(2m-1)^{2}}{4}t-\frac{r^{2}}{4t}}t^{-\frac{5+sp}{2}}\,\mathrm{d}t&=2(2m-1)^{\frac{3+sp}{2}}\widetilde{K}_{\frac{3+sp}{2},\frac{2m-1}{2}}(r)\\ &=4(2m-1)^{\frac{1+sp}{2}}\left(\frac{-\partial_{r}}{r}\right)\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r).\end{split}$
@@
$\begin{split}&C_{2}\int_{0}^{\infty}e^{t\Delta_{\mathbb{H}^{n}}}[g_{\varepsilon}(x,\cdot)](x)\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}\\ &=c_{n,s,p}C_{1}\int_{\mathbb{H}^{n}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{(-\partial_{r})\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r\,g_{\varepsilon}(x,\xi)\,\mathrm{d}\xi\\ &=c_{n,s,p}C_{1}\int_{\mathbb{H}^{n}}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,g_{\varepsilon}(x,\xi)\,\mathrm{d}\xi,\end{split}$
where we used Lemma 3.1 (m1)(m-1)-times with $\nu=\frac{1+sp}{2}$, $a=\frac{2m-1}{2}$, and $y=0$ in the last equality. This proves (3.5) in the even-dimensional case.